OpenStudy (trojanpoem):

ABC is a triangle such that BC * AD * AB = 200 cm^3 The area of the triangle is 40 cm^2 then the radius of the circle passing through it's heads is ... cm

OpenStudy (welshfella):

what is AD?

OpenStudy (trojanpoem):

AC sorry typo

OpenStudy (phi):

circle passing through it's heads does this passing through the vertices A,B, and C ?

OpenStudy (trojanpoem):



OpenStudy (phi):

it sounds hard.

OpenStudy (phi):

any hints? what are you studying ?

OpenStudy (trojanpoem):

5 hours thinking and no result :( The question is in general questions , I also don't know the lesson name.

OpenStudy (phi):

general questions?

OpenStudy (trojanpoem):

I forgot  The question was in determinators although I don't know what's the relation.

OpenStudy (trojanpoem):

The picture = 200 cm^3

OpenStudy (phi):

wikipedia says diameter of the circumcircle = abc/2*area so the radius is abc/area

OpenStudy (welshfella):

how do you find the area?

OpenStudy (trojanpoem):

It's given , but what is abc ?

OpenStudy (welshfella):

of course silly me abc = 200

OpenStudy (welshfella):

a, b and c are the 3 sides of the traingle

OpenStudy (trojanpoem):

200/40 = 5 ? the book says the final answer is 5/4

OpenStudy (welshfella):

os that correct for the radius @phi ?

OpenStudy (phi):

I was looking at this https://en.wikipedia.org/wiki/Circumscribed_circle

OpenStudy (trojanpoem):

is 5/4 logical ?

OpenStudy (welshfella):

yes I think it is radius = abc / (4*area)

OpenStudy (phi):

oops! what's a multiple of 2 one way or the other!

OpenStudy (phi):

diameter of the circumcircle = abc/2*area so radius is abc/4*area

OpenStudy (welshfella):

yea lol

OpenStudy (trojanpoem):

Yeah.

OpenStudy (trojanpoem):

Thanks so much guys.

OpenStudy (welshfella):

yw

OpenStudy (phi):

but it's not obvious why it works

OpenStudy (trojanpoem):

Where can I find how did they prove it ?

OpenStudy (welshfella):

yea it's quite obscure

OpenStudy (phi):

I would google.

OpenStudy (phi):

looks promising

OpenStudy (welshfella):

yea